OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = (4(4n+1)(8n +7))/(6)

OpenStudy (raden):

put n=1 to LHS and RHS LHS : 4*1(4*1+2) = 4 * 6 = 24 RHS : 4(4*1+1)(8*1 +7))/(6) = 4(5)(15)/6 = 50 because LHS not equal RHS, so it is false

OpenStudy (anonymous):

thanks!

OpenStudy (raden):

welcome